OpenStudy (anonymous):

(72 + 5 * 88 + 72 + 93 - 53 + 78 * 3)/5

OpenStudy (anonymous):

\[\frac{1}{5} (72+5\ 88+72+93-53+78\ 3)=\frac{858}{5}=171.6 \]